40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 24 02:12:05 2014, Jenne, Update, LSC, Locking without TRY TRX_avg_40cts_POPDC_avg_200cts.pngLockloss_PRCL_350Hz.pngLockloss_DARM_20Hz.png
    Reply  Thu Apr 24 11:19:50 2014, Koji, Update, LSC, Locking without TRY 
       Reply  Thu Apr 24 14:00:42 2014, Jenne, Update, LSC, Locking without TRY REFLDC_1sec.pngREFLDC_3peaks.pngREFLDC_1peak.pngREFLDC_1peak_zoom.png
          Reply  Mon Apr 28 13:18:18 2014, Jenne, Update, LSC, LSC offsets script modified 
Message ID: 9868     Entry time: Mon Apr 28 13:18:18 2014     In reply to: 9848
Author: Jenne 
Type: Update 
Category: LSC 
Subject: LSC offsets script modified 

Quote:

The weird jumps at the beginning of each TRX peak are due to the triggered switching between the Thorlabs trans PD and the QPD trans PD.  Clearly we need to work on their relative normalizations.  There are also little jumps after each peak as the triggering sends the signal back to the Thorlabs PD.

 I was unhappy with the discontinuities between the Thorlabs and QPD versions of our transmitted light powers.  I realized that in the olden days, we just used the Thorlabs PD, and we set the no-light offset in the LSC version of the TR[x,y] filter banks.  However, now that we have brought the QPDs back, we are setting the dark offsets in the end suspension models, so that the signal chosen by the trigger already has its offset taken care of before we send it to the LSC model. 

Anyhow, having the offsets script try to put a value in the C1:LSC-TR[x,y]_OFFSET was giving us an extra offset and then when we did the normalizations, the numbers came out all wrong.  So.  I have removed the C1:LSC-TR[x,y] filter banks from the offset list, since they were made redundant. 

I have redone the normalizations for both arms (after running the ASS scripts).  I checked by watching the _OUT16 versions of the Thorlabs and QPD diodes before the triggering happens, and as I put offsets into the LSC servos to change the transmitted power, the diodes both change in the same way.  So, we'll have to see if this holds true for more than just values 0-1 the next time we lock, but hopefully it won't need changing for a while.

ELOG V3.1.3-